An incompressible fluid flows through a horizontal pipe. At one point in the pipe the pressure in the fluid is p1 and the fluid speed is v1. further down the pipe the pressure is p2 and the speed is 2v1. what can be concluded about p1 and p2?

Grade:12th pass

1 Answers

Vikas TU
7 years ago
From Bernoulli’s principle,
p1 + 0.5p*v1^2 = p2 + 0.5p(2v2)^2
we get,
p1  +  p*v1^2/2 = p2 + 2pv2^2
or
p1 – p2 = 2*p*v2^2 – p*v1^2/2
This would be the relation.